itsnaenae:

A paint can has a diameter of 17 centimeters and a height of 24 centimeters. How many cubic centimeters of paint will fill the can? 5,444.76 cm3 21,779.04 cm3 1,281.12 cm3 7,686.72 cm3

Narad:

Volume of paint is \[= \pi *(d ^{2}/4)*h\] \[= \pi * 17^{2}/4*24\] =\[=5444.76 cm ^{3}\]